OpenStudy (anonymous):

HELP !?!?! A point H on a segment with endpoints B (3, −1) and Z (12, 5) partitions the segment in a 5:1 ratio. Find H. You must show all work to receive credit.

OpenStudy (anonymous):

@Directrix should i use the midpoint formula ??

OpenStudy (rajat97):

i think you should use the internal division formula take a look at this link this is good http://www.teacherschoice.com.au/Maths_Library/Analytical%20Geometry/AnalGeom_3.htm

OpenStudy (anonymous):

how to use it ??

Directrix (directrix):

The midpoint formula is not going to help here as best I see. @luka1998 Did you calculate the length of segment ZB? If so, what is it?

OpenStudy (anonymous):

BY using Distance right ?

Directrix (directrix):

Yes.

Directrix (directrix):

@rajat97 Feel free to show us how to use this formula and apply it to this problem.

OpenStudy (rajat97):

yup i'm on it @Directrix :)

OpenStudy (rajat97):

let the point B(3,-1) be (x1,y1) and let the point Z (12,5) be (x2,y2) and let the segment be divided by the line in the ratio of 5:1 that is m:n by the point (x,y) so by the internal division formula, x=\[(n \times x1 + m \times x2)/m+n\] and for the y-co-ordinate, y=\[(n \times y1 + m \times y2)/m+n\]

OpenStudy (anonymous):

@Directrix d=squar root (x2-x1)^2+(y2-y1)^2 ZB=(12,5) (3,-1) squar root (3-12)^2+(-1-5)^2 squar root 81+36 = squar root 117

OpenStudy (rajat97):

i'll explain this in a better way but for this question , you can use what i have written just wait for a minute

OpenStudy (rajat97):

 let q be the point that divides the line seg. PR in the ratio PQ:QR=m:n so you should use the formula like this m is the part of the ratio , farthest from the point R so m goes in multiplication with the coordinates of point R for finding the co-ordinates and the same goes with n and it goes in multiplication with the co-ordinates of P if you don't understand this thing, just look at the formula that i have posted in my previous post and then read this once again carefully this is ofcourse not the proof for this formula (lol) this is just for remembering purpose hope this helps you !

OpenStudy (anonymous):

@rajat97 is this correct B(3,−1) Z(12,5) ratio of 5:1 at point P(x,y) P(x,y) ( 5(3)+1(12)over 5+1, 5(−1)+1(5)over 5+1 ) (27 over 6, 0 over 6 ) (4.5 , 0)

OpenStudy (rajat97):

nope! you've done it reverse i think it should be B(3,−1) Z(12,5) ratio of 5:1 at point P(x,y) P(x,y) ( 1(3)+5(12)over 5+1, 1(−1)+5(5)over 5+1 ) so this gives you (63over6 , 24 over 6) so p is (10.5 , 4) hope this is correct! well try!

OpenStudy (rajat97):

and thanks for the medal

OpenStudy (anonymous):

you are welcome and thx

OpenStudy (rajat97):

it's my pleasure to help you

OpenStudy (rajat97):

Thanks for the medals everybody!